1. You diagnosed 7-year-old Emma with right otitis media. Emma weighs 81 lbs. You order amoxicillin 90 mg/kg/day divided every 8 hours for seven days. Amoxicillin chewable comes in 125 mg or 250 mg strength.

a) What is Emma’s total daily dose?


b) What is Emma’s individual daily dose?


c) What strength are you ordering and how many chewables will you order?
Answer #1

Ans) Weight: 81 lb = 81 ÷ 2.205 = 36.74 kg

Order: 90 mg/kg/day = 90 × 36.74 = 3306.6 mg.

Every 8 hours means 3 times in a day.

Available: Amoxillin 125 mg or 250 mg

a) Daily dose: 3306.6 mg

b) Individual daily dose: 1102.2 mg

c) Amoxillin chewables for Individual dose will be: 3 tablets of 125 mg & 3 tablets of 250 mg.